# Tracing Intricate Cues in Dialogue: Joint Graph Structure and Sentiment Dynamics for Multimodal Emotion Recognition

### v1

@@ V-D Time and Memory Overhead
In Table VII, we report the running time and memory usage of our model and its baselines for the MERC task across all datasets. Relative to other models, MMGCNN and GraphSmile exhibit an absolute advantage in time expenditure on the IEMOCAP dataset, yet MMGCN’s time usage on the other two datasets is significantly higher than that of GraphSmile. M3Net and GraphSmile achieve lower time consumption on the MELD dataset, while M3Net’s running time on other datasets is inferior to GraphSmile. The running times of DialogueCRN, MM-DFN, and SACL-LSTM across all datasets are unsatisfactory. MMGCN’s memory consumption is considerably lower than other models on all datasets, but its overly simplistic model leads to poor performance on most datasets. M3Net and SACL-LSTM have much higher memory consumption on multiple datasets compared to other models, and the memory expenditure of DialogueCRN and MM-DFN is also higher than that of the proposed model. In summary, GraphSmile not only achieves the best accuracy and F1 scores but also balances the costs of time and memory.
@@ VI Conclusion
In future work, we intend to address other salient challenges in the realm of affective computing in conversation. For instance, tackling the differentiation of similar emotions and issues of class imbalance through contrastive learning strategies. We will also explore more sophisticated multimodal fusion strategies to capitalize on the unique strengths of each modality and enhance the model’s robustness against noise and irrelevant information.

### v2

@@ V-D Time and Memory Overhead
In Table VII, we report the running time and memory usage of our model and its baselines for the MERC task across all datasets. Note that all these experiments are conducted on a single NVIDIA GeForce RTX 3090. Relative to other models, MMGCNN and GraphSmile exhibit an absolute advantage in time expenditure on the IEMOCAP dataset, yet MMGCN’s time usage on the other two datasets is significantly higher than that of GraphSmile. M3Net and GraphSmile achieve lower time consumption on the MELD dataset, while M3Net’s running time on other datasets is inferior to GraphSmile. The running times of DialogueCRN, MM-DFN, and SACL-LSTM across all datasets are unsatisfactory. MMGCN’s memory consumption is considerably lower than other models on all datasets, but its overly simplistic model leads to poor performance on most datasets. M3Net and SACL-LSTM have much higher memory consumption on multiple datasets compared to other models, and the memory expenditure of DialogueCRN and MM-DFN is also higher than that of the proposed model. In summary, GraphSmile not only achieves the best accuracy and F1 scores but also balances the costs of time and memory.
@@ VI Conclusion
In future work, we intend to address other salient challenges in the realm of affective computing in conversation. For instance, (1) tackling the differentiation of similar emotions and issues of class imbalance through contrastive learning strategies, and (2) delving deeper into the sentimental dynamics in conversation by considering the temporal logic of transitions and refining the types of shift labels. We will also explore more sophisticated multimodal fusion strategies to capitalize on the unique strengths of each modality and enhance the model’s robustness against noise and irrelevant information. Recently, large language models (LLMs) have demonstrated formidable reasoning capabilities, thus integrating MERC and MSAC with LLM technologies will further propel the development of affective community.